OpenStudy (anonymous):

Flaws in a carpet ten to occur randomly and independently at a rate of two every 100 square feet. What is the probability that: a) a carpet that is 8 feet by 12 feet contains three flaws

OpenStudy (horsegirl27):

Well, what do you think?

OpenStudy (horsegirl27):

Start by finding the area of the carpet

OpenStudy (horsegirl27):

A=l*w length=8 width=12 A=8*12 A=?

OpenStudy (anonymous):

A=lxw =8x15 =120

OpenStudy (anonymous):

@horsegirl27

OpenStudy (horsegirl27):

Oops, you had 8*15, you want 8*12

OpenStudy (anonymous):

Oh you know what in the question i typed 12 its supposed to be 15 sorry for the confusion. But now that I have the area then what? @horsegirl27

OpenStudy (horsegirl27):

So, if a carpet will have 2 flaws every 100 sq ft, and the area is 96 sq ft, is there a chance there will be 3 flaws?

OpenStudy (anonymous):

No I guess not because there is not enough area but if the area is 120 which is the 8*15 then it would have enough room to have 3 flaws right? @horsegirl27

OpenStudy (horsegirl27):

I'm not sure. But since your area is 96 sq ft, your answer will be no.

OpenStudy (horsegirl27):

Well, not exactly no. Next you will have to find the probability

OpenStudy (anonymous):

If you look at my messages above it says that I made a mistake and that the 12 should be a 15 @horsegirl27

OpenStudy (horsegirl27):

oh right, I'm sorry I thought you meant when you typed it in the problem.

OpenStudy (anonymous):

its supposed to be 8*15 so how do I solve this? @horsegirl27

OpenStudy (mathmate):

@Shaekitchen hint: review Poisson distribution in your notes. \(\lambda\) equals to 2*96/100.

OpenStudy (anonymous):

Is that my full answer thought? Im not sure where else to take it? My prof is crap at writing notes I cant understand it.@mathmate

OpenStudy (mathmate):

What I am saying is the probability is obtained by the Poisson distribution, with lambda = 0.96*2/100= expected number of flaws for 96 sq.ft, and n=3 (3 flaws).

OpenStudy (mathmate):

You have to do a calculation to find P(n)=\(\large \frac{\lambda ^k}{k!}e^{-\lambda}\) \(\lambda = 2*96/100 =expected number of flaws, k=3 (flaws) ) For further information, read: http://en.wikipedia.org/wiki/Poisson_distribution

OpenStudy (mathmate):

* P(k) = ....